OpenStudy (anonymous):

Factor Completely 21x^2 + 31x + 4

OpenStudy (anonymous):

lotta choices

OpenStudy (anonymous):

( 7 + 2 ) ( 2+ 2 )

OpenStudy (anonymous):

i would go with \((7x+...)(3x+...)\) and see what words

OpenStudy (anonymous):

is that the answer ? factoring what ?

OpenStudy (anonymous):

this is a pain two numbers whose product is 4, could be 1 and 4 or 2 and 2, but you need the sum to be 31 which is odd, so 2 and 2 is out find the right place to put the 1 and 4

OpenStudy (anonymous):

answer is easy enough to come by, but trial and error is the method

OpenStudy (callisto):

21x^2 + 31x + 4 = 21x^2 + 3x + 28x + 4 = 3x ( 7x +1) + 4 (7x+1) = (3x+4)(7x+1)